OpenStudy (anonymous):

jerome is five years older than gerald .three years ago he was twice as old as gerald how old are they?

OpenStudy (anonymous):

J is Jerome G is Gerald J=G+5 J-3=2(G-3) J=2G-3 Substitution: G+5=2G-3 G=8 J=8+5=13 Jerome is 13, Gerald is 8